The Spartan National Bank has agreed to lend you $30,000 today. In three years you must pay them $42,147. What is the annual rate of interest?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

42147=30,000*(1+r/100)^3

(42147/30,000)^(1/3)=(1+r/100)

(1+r/100)=1.12

r=1.12-1

=12%(Approx).